### v1

@@ 3 Model
The second term and third terms of the utility function (1) are externalities generated while linking to package $j$. Indeed, package $j$ may be linked to other packages $r$ as well. This means that the coder needs to check the quality and features of the code in these packages, to make sure they are free of bugs and compatible with the coder’s own code. On the other hand, any update to package $r$ may compromise compatibility to package $j$ and $i$, so this also includes costs of maintaining. In short, $w_{ijr}(\gamma)$ measures the net benefits of this externality when coders of $i$ form a dependency to library $j$.
Furthermore, if package $i$ creates a dependencies to library $j$ it may compromise compatibility with other packages $r$ that have a dependencies to $i$. This is accounted for by the third term $v_{ijr}(\gamma)$ in utility function (1).
We assume that the second and third term in the utility take the form
$w_{ijr}(\gamma)=v_{ijr}(\gamma)=\begin{cases}\gamma&\text{ if }z_{i}=z_{j}=z_{r}\\ 0&\text{otherwise,}\end{cases}$ (3)
This specification assumes that the externalities are local and have the same functional form. The reason for this normalization is that it guarantees the existence of a potential function that characterizes the equilibrium. Furthermore, this assumptions facilitates identification for parameter $\gamma$, through variation within types (Mele, 2017, 2022; Schweinberger and Handcock, 2015; Schweinberger and Stewart, 2020; Babkin et al., 2020).
To summarize, the coders need to make decisions on whether to link another package or not. When making the decision, they take into account some externalities from this network of dependencies, but not all the possible externalities.
We assume a dynamic process of package development. Time is discrete and in each time period $t$ only one coder is making decisions. At time $t$, a randomly chosen package $i$ needs some code update. Package $j$ is proposed for the update. If package $j$ is not already linked to package $i$, we have $g_{ij,t-1}=0$. If the coder decides to write the code in-house there is no update and $g_{ij,t}=0$. On the other hand, if the coder decides to link to package $j$, the network is updated and $g_{ij,t}=1$. If the dependency already exists (i.e. $g_{ij,t-1}=1$), then the coder decision is whether to keep the dependency ($g_{ij,t}=1$) or unlink package $j$ and write the code in-house instead ($g_{ij,t}=0$). Here, we make the simplifying assumption that coders cannot substitute a dependency for another dependency at the same time. Formally, with probability $\rho_{ij}>0$ package $i$ coders propose an update that links package $j$. The coder decides whether to create this dependency or write the code in-house. Before linking the coder receives a random shock to payoffs $\varepsilon_{ij0},\varepsilon_{ij1}$, which is i.i.d. among packages and time periods. This shock captures that unexpected things can happen both when developing code in house and when linking to an existing package.
So the coders of $i$ form a dependency link to package $j$ if:
$U_{i}\left(\bm{g^{\prime}},\bm{x},\bm{z};\bm{\theta}\right)+\varepsilon_{ij1}\geq U_{i}\left(\bm{g},\bm{x},\bm{z};\bm{\theta}\right)+\varepsilon_{ij0}$ (4)
where $\bm{g^{\prime}}$ is network $\bm{g}$ with the addition of link $g_{ij}$.
Throughout the paper we maintain the following assumptions (Mele, 2017):
  1. 1.
    The probability that coders of $i$ propose an update that links to package $j$ is strictly positive, $\rho_{ij}>0$ for all pairs $i,j$. This guarantees that any dependencies can be considered.
  2. 2.
    The random shock to payoffs, $\varepsilon_{ij}$, follows a logistic distribution. This is a standard assumption in many random utility models for discrete choice.
As shown in Mele (2017) and Mele (2022), after conditioning on the unobservable types $\bm{z}$, the sequence of networks generated by this process is a Markov Chain and it converges to a unique stationary distribution that can be expressed in closed-form as a discrete exponential family with normalizing constant. In our model, the stationary distribution is
@@
$\pi(\bm{g},\bm{x},\bm{z};\bm{\theta})$$\coloneqq$ $\prod_{k=1}^{K}\frac{e^{Q_{kk}}}{c_{kk}}\left[\prod_{l\neq k}^{K}\prod_{i=1}^{N_{k}}\prod_{j=1}^{N_{l}}\frac{e^{u_{ij}(\alpha_{b},\bm{\beta}_{b})}}{1+e^{u_{ij}(\alpha_{b},\bm{\beta}_{b})}}\right],$ (5)
where the potential function $Q_{kk}$ can be written as:
$Q_{kk}(\bm{g}_{kk})$ ::=:= $Q\left(\bm{g}_{kk},\bm{x},\bm{z};\alpha_{w},\bm{\beta}_{w}\right)$ (6)
@@
$=$ $\sum_{i=1}^{N}\sum_{j\neq i}^{N}g_{ij}z_{ik}z_{jk}\left(\alpha_{w}+\sum_{p=1}^{P}\beta_{wp}\bm{1}\{\bm{x}_{ip}=\bm{x}_{jp}\}\right)+\gamma\sum_{i=1}^{N}\sum_{j\neq i}^{N}\sum_{r\neq i,j}^{N}g_{ij}g_{jr}z_{ik}z_{jk}z_{rk},$
and $\bm{g}_{kk}$ is the network among nodes of type $k$; the normalizing constant is:
$c_{kk}\;\coloneqq\;\sum_{\omega_{kk}\in\mathcal{G}_{kk}}e^{Q_{kk}(\bm{\omega}_{kk})}.$ (7)
Here, $\bm{\omega}_{kk}\in\mathcal{G}_{kk}$ is one network in the set of all possible networks among nodes of type $k$ [^20] As noted in Mele (2017), the constant $c_{kk}$ is a sum over all $2^{N_{k}(N_{k}-1)/2}$ possible network configurations for the nodes of type $k$. This makes the computation of the constant impractical or infeasible in large networks.
The stationary distribution $\pi$ represents the long-run distribution of the network, after conditioning on the unobservable types $\bm{z}$. The second product in (5) represents the likelihood of links between packages of different unobservable types, while the first part is the likelihood of links among libraries of the same type. This simple decomposition is possible because the externalities have been normalized to be local. So the model converges to $K$ independent exponential random graphs for links of the same type, and implies conditionally independent links for between-types connections.
Because of this simple characterization of the stationary equilibrium, the incentives of the coders are summarized by a potential function whose maxima are pairwise stable networks. Because of the externalities, in general, the equilibrium networks will be inefficient, as they will not maximize the sum of utilities of the coders. [^21] This is evident when we compute the sum of utilities. Indeed, the welfare function will include all the externalities and, therefore, will be, in general, different from the potential, where the externalities are multiplied by 1/2. See also Mele (2017) for more details.
@@ 4.1 Estimation
Our novel method for estimation of a directed network formation model with observable and unobservable types is a crucial contribution of this paper. Although scalable estimation methods are available for simpler models, like stochastic blockmodels (Vu et al., 2013), conditionally independent links models (Mele et al., 2023), or models for undirected networks (Dahbura et al., 2021; Babkin et al., 2020; Schweinberger and Stewart, 2020), there is no scalable method for directed networks readily available.
To understand the computational problem involved in this estimation, we write the likelihood of our model as
$\mathcal{L}(\bm{g},\bm{x};\bm{\alpha},\bm{\beta},\gamma,\bm{\eta})=\sum_{\bm{z}\in\mathcal{Z}}L\left(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\gamma,\bm{\eta}\right)=\sum_{\bm{z}\in\mathcal{Z}}p_{\bm{\eta}}\left(\bm{z}\right)\,\pi(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\gamma).$ (8)
where each node’s type distribution $p_{\bm{\eta}}(\bm{z})$ is multinomial, and types are i.i.d.
[…]
Maximizing (8) involves a sum over all possible type allocations of each node. Furthermore, the stationary distribution $\pi(\bm{g},\bm{x},\bm{z};\bm{\alpha},\bm{\beta},\gamma)$ (the conditional likelihood) is a function of an an intractable normalizing constant. Therefore direct computation of the objective function is infeasible.
Our strategy is to separate the estimation of unobserved types from the structural utility parameters according to a two-step algorithm. First, we estimate the type of each node and, second, we estimate the structural parameters $\bm{\alpha},\bm{\beta},$ and $\gamma$, conditional on the type of each node (Babkin et al., 2020).
In the first step, we extend variational approximations with maximization minorization updates to directed dependencies for approximating the intractable likelihood and estimate the posterior distribution of types (Vu et al., 2013). We then assign the types to each node according to the highest posterior probability. This step exploits the fact that our model corresponds to a stochastic blockmodel with covariates when $\gamma=0$, i.e. when there are no link externalities. Although the variational mean-field approximation algorithm leads to a sequential algorithm, we augment a lower bound of the likelihood by maximizing a minorizing function, which is quadratic. Thereby, we have to solve a quadratic maximization problem whose terms are calculated via fast sparse matrix operations and whose solution is found via quadratic problem solvers with box constraints (Stefanov, 2004). Through this approximation, we can perform the estimation of unobserved types at scale while preserving the complex correlation of dependency links in our data (more details are provided in Appendix A).
In the second step, we estimate the remaining structural parameters using a pseudolikelihood estimator that maximizes the conditional choice probabilities of links conditional on the estimated types (Schweinberger and Stewart, 2020; Dahbura et al., 2021). This is a computationally feasible estimator that scales well with the size of the network. We provide the theoretical and computational details on the scalable estimation procedure in Appendix A.

### v2

@@ 3.1 Setup
The second and third term of utility function (2) are externalities generated by linking to package $j$, because package $j$ may be linked to other packages $r$ as well. This means that the developers of package $i$ need to check the quality and features of the code in those packages $r$ as well, to ensure that those packages are free of bugs or vulnerabilities, and compatible with package $i$. On the other hand, any update of package $r$ may compromise compatibility with packages $i$ and $j$, increasing the cost of maintaining package $i$. In short, $w_{ijr}$ measures the net benefits of this externality when developers of package $i$ form a dependency to package $j$. In addition, if the developers of package $i$ create a dependency on package $j$, it may compromise compatibility with other packages $r$ that depend on package $i$. This is accounted for by the third term $v_{ijr}$ in utility function (2). We assume that the second and third term in the utility are of the form
$v_{ijr}(\gamma)\;=\;w_{ijr}(\gamma)\;\coloneqq\;\begin{cases}\gamma&\text{ if $i$, $j$, $r$ are all of type $k$}\\ 0&\text{otherwise,}\end{cases}$ (6)
where $\gamma\in\mathbb{R}$.
The choice of the utility functions $U_{i}$ guarantees that all externalities are local, in the sense that all externalities are limited to packages of the same type. Local externalities make sense in large networks, because developers are unlikely to know the functionalities of thousands of other packages and, more likely than not, will create dependencies to other packages of the same type, which developers are most familiar with. In addition, local externalities have probabilistic, statistical, and computational advantages. First, local externalities ensure desirable properties of models for large networks (Schweinberger and Handcock, 2015; Schweinberger et al., 2020). Second, local externalities facilitate theoretical guarantees for statistical procedures (Schweinberger and Stewart, 2020; Schweinberger, 2020; Stewart and Schweinberger, 2023; Stewart, 2024). Third, local externalities enable local computing on subnetworks of large networks, which facilitates parallel computing on a massive scale (Babkin et al., 2020).
3.2 Equilibrium
Software development is a dynamic process. We assume that time is discrete and that at each time point $t=1,2,\ldots$ a single package is updated. In other words, maintainers do not update software packages continuously, but certain events – unobservable to researchers – trigger the need for an update. For example, if package $i$ depends on package $j$, then an update of package $j$ may require an update of package $i$, or a request by users of package $i$ may trigger an update of package $i$. The update can be produced in-house or by forming a dependency to some other package $j$ containing the same functionality. Since researchers may not be able to observe the timing of such updates, we model updates as a discrete-time Markov process (Norris, 1997).
The first ingredient if the discrete-time Markov process is the update process that determines which package $i$ requires an update and which dependency $j$ is considered to implement the update. We assume that the probability that package $i$ requires an update and the developers of package $i$ consider creating, maintaining, or removing a dependence to package $j$ is $\rho(\bm{g}^{t-1},\bm{x}_{i},\bm{x}_{j},\bm{z}_{i},\bm{z}_{j})$, where $\bm{g}^{t-1}$ is the network at time $t-1$. So the update process may depend on the existing network $\bm{g}^{t-1}$ as well as observables ($\bm{x}_{i}$, $\bm{x}_{j}$) and unobservables ($\bm{z}_{i}$, $\bm{z}_{j}$). A simple example is an update process that selects an ordered pair of packages $i$ and $j$ at random. In other words, a package $i$ is selected with probability $1/n$ and requires an update. Conditional on the event that package $i$ requires an update, the developers of package $i$ consider updating the possible dependency of package $i$ on package $j$ with probability $1/(n-1)$. As a result, the probability that package $i$ requires an update and the developers of package $i$ consider updating a (possible) dependency of package $i$ on package $j$ is
$\rho\left(\bm{g}^{t-1},\bm{x}_{i},\bm{x}_{j},\bm{z}_{i},\bm{z}_{j}\right)\;\;=\;\;\dfrac{1}{n(n-1)}.$
A second example is an update process that allows the conditional probability that an update of the (possible) dependency of package $i$ on package $j$ is proposed to depend on the distance $d(x_{i},x_{j})$ between packages $i$ and $j$:
$\rho\left(\bm{g}^{t-1},\bm{x}_{i},\bm{x}_{j},\bm{z}_{i},\bm{z}_{j}\right)\;\;\propto\;\;\frac{1}{n}\exp\left(-d(\bm{x}_{i},\bm{x}_{j})\right),$
where $d(\bm{x}_{i},\bm{x}_{j})$ is a function of observables ($\bm{x}_{i}$, $\bm{x}_{j}$) satisfying reflexity, symmetry, and the triangle inequality. In other words, $d(\bm{x}_{i},\bm{x}_{j})$ is a distance function that quantifies the distance between packages $i$ and $j$ based on observable characteristics $\bm{x}_{i}$ and $\bm{x}_{j}$ of packages $i$ and $j$.
[…]
$\rho\Big{(}\bm{g}^{t-1},\bm{x}_{i},\bm{x}_{j},\bm{z}_{i},\bm{z}_{j}\Big{)}\;\;=\;\;\rho\Big{(}\bm{g}_{-(i,j)}^{t-1},\bm{x}_{i},\bm{x}_{j},\bm{z}_{i},\bm{z}_{j}\Big{)}.$
Second, the update probabilities $\rho(\bm{g}_{-(i,j)}^{t-1},\bm{x}_{i},\bm{x}_{j},\bm{z}_{i},\bm{z}_{j})$ are strictly positive for all ordered pairs of packages $(i,j)$:
[…]
This implies that each possible dependency can be proposed with a positive probability $\rho$, no matter how small. These assumptions help ensure that in the long-run equilibrium these probabilities do not affect the distribution of networks (Mele, 2017).
While the update process is stochastic, the decision to create, maintain, or remove a possible dependency of package $i$ on package $j$ is a strategic decision. Conditional on the opportunity to update the dependency of package $i$ on package $j$, the developers of package $i$ make a decision about whether to create, maintain, or remove a dependency on package $j$ based on utility function $U_{i}(\bm{g};\,\bm{x},\bm{z},\bm{\theta})$. Before the developers of package $i$ make a decision about the (possible) dependency on package $j$, the developers receive random shocks $\varepsilon_{ij0}\in\mathbb{R}$ and $\varepsilon_{ij1}\in\mathbb{R}$ to the perceived utility of a dependency of package $i$ on package $j$ ($g_{ij}^{t}=1$) and the perceived utility of writing code in-house ($g_{ij}^{t}=0$). After receiving random shocks $\varepsilon_{ij0}\in\mathbb{R}$ and $\varepsilon_{ij1}\in\mathbb{R}$, the developers of package $i$ create a dependency on package $j$ if the perceived utility of the dependency of package $i$ on package $j$ ($g_{ij}^{t}=1$) exceeds the utility of writing code in-house ($g_{ij}^{t}=0$):
$U_{i}\left(g_{ij}^{t}=1,\,\bm{g}_{-(i,j)}^{t-1};\,\bm{x},\bm{z},\bm{\theta}\right)+\varepsilon_{ij1}^{t}\;\geq\;U_{i}\left(g_{ij}^{t}=0,\,\bm{g}_{-(i,j)}^{t-1};\,\bm{x},\bm{z},\bm{\theta}\right)+\varepsilon_{ij0}^{t},$ (7)
where $\bm{g}_{-(i,j)}^{t-1}$ refers to network $\bm{g}^{t-1}$ excluding the possible dependency $g_{ij}$ of package $i$ on package $j$. The random shocks to payoffs $\varepsilon_{ij0}^{t}\in\mathbb{R}$ and $\varepsilon_{ij1}^{t}\in\mathbb{R}$ capture the fact that unexpected events can occur when developing code in-house ($\varepsilon_{ij0}$) and when creating a dependency ($\varepsilon_{ij1}$), e.g., changes in the composition of the developer team or additional information about package $j$. We assume that the random shocks to payoffs $\varepsilon_{ij0}$ and $\varepsilon_{ij1}$ do not depend on time, are independent across pairs of packages $(i,j)$, and follow a logistic distribution, which is a standard assumption in the literature on discrete choice models (Mele, 2017; Graham and dePaula, 2020; Boucher and Mourifie, 2017). As a consequence, when the opportunity to update the dependency of package $i$ on package $j$ arises, the conditional choice probability of a dependency of package $i$ on package $j$ is
$\mathbb{P}\left(g_{ij}^{t}=1\ |\ \bm{g}_{-(i,j)}^{t-1};\,\bm{x},\bm{z},\bm{\theta}\right)=\frac{\exp\left(U_{i}(g_{ij}^{t}=1,\,\bm{g}_{-(i,j)}^{t-1};\,\bm{x},\bm{z},\bm{\theta})-U_{i}(g_{ij}^{t}=0,\,\bm{g}_{-(i,j)}^{t-1};\,\bm{x},\bm{z},\bm{\theta})\right)}{1+\exp\left(U_{i}(g_{ij}^{t}=1,\,\bm{g}_{-(i,j)}^{t-1};\,\bm{x},\bm{z},\bm{\theta})-U_{i}(g_{ij}^{t}=0,\,\bm{g}_{-(i,j)}^{t-1};\,\bm{x},\bm{z},\bm{\theta})\right)}.$ (8)
As shown in Mele (2022) (with unobservables) and Butts (2009) and Mele (2017) (without unobservables), the sequence of networks generated by the process described above is a discrete-time Markov process (Norris, 1997), conditional on the unobservable types $\bm{z}$.
The discrete-time Markov chain described above is irreducible and aperiodic and hence admits a unique limiting distribution $\pi$, which is the stationary distribution of the Markov chain (Norris, 1997). The stationary distribution $\pi$ can be represented in closed form as an exponential-family distribution (Mele, 2017) and is given by
@@ 3.2 Equilibrium
$\pi(\bm{g};\,\bm{x},\,\bm{z},\,\bm{\theta})$$\coloneqq$ $\left(\prod_{k=1}^{K}\frac{e^{Q_{kk}(\bm{g}_{kk};\,\bm{x},\,\bm{z},\,\bm{\theta})}}{c_{kk}(\bm{x},\,\bm{z};\,\bm{\theta})}\right)\;\left(\prod_{l\neq k}^{K}\,\prod_{i=1}^{N_{k}}\,\prod_{j=1}^{N_{l}}\frac{e^{u_{ij}(\bm{\alpha}_{b},\,\bm{\beta}_{b})}}{1+e^{u_{ij}(\bm{\alpha}_{b},\,\bm{\beta}_{b})}}\right).$ (9)
The potential function $Q_{kk}$ of the within-type-$k$ distribution is defined as
@@
$\begin{array}[]{lllll}Q_{kk}\left(\bm{g}_{kk};\,\bm{x},\,\bm{z},\,\bm{\theta}\right)&\coloneqq&\displaystyle\sum_{i=1}^{N}\,\sum_{j\neq i}^{N}u_{ij}\left(\bm{\alpha}_{w},\bm{\beta}_{w},\bm{x_{i}},\bm{x}_{j},\bm{z}_{i},\bm{z}_{j}\right)g_{ij}z_{ik}z_{jk}\vspace{.25cm}\\ &+&\gamma\,\displaystyle\sum_{i=1}^{N}\,\sum_{j\neq i}^{N}\,\sum_{r\neq i,j}^{N}g_{ij}g_{jr}z_{ik}z_{jk}z_{rk},\end{array}$ (10)
while the normalizing constant $c_{kk}(\bm{x},\,\bm{z};\,\bm{\theta})$ of the within-type-$k$ distribution is defined as
$c_{kk}(\bm{x},\,\bm{z},\,\bm{\theta})$$\coloneqq$ $\displaystyle\sum_{\bm{\omega}\,\in\,\mathcal{G}_{kk}}e^{Q_{kk}(\bm{\omega};\;\bm{x},\,\bm{z},\,\bm{\theta})},$ (11)
where the sum is over all possible networks $g_{kk}^{\prime}\in\mathcal{G}_{kk}$ of type $k$.
The limiting distribution $\pi$ of the discrete-time Markov chain represents the long-run distribution of the software development process, conditional on the types $\bm{z}$ of packages. The first term in (9) represents the likelihood of dependencies among packages of the same type, while the second term represents the likelihood of dependencies between packages of different types.
The stationary distribution $\pi$ implies that the incentives of developers are captured by a potential function of the form
$Q(\bm{g},\bm{x},\bm{z},\bm{\theta})$ $=$ $\sum_{k=1}^{K}Q_{kk}\left(\bm{g}_{kk};\,\bm{x},\,\bm{z},\,\bm{\theta}\right)+\sum_{k=1}^{K}\sum_{l\neq k}Q_{kl}\left(\bm{g}_{kl};\,\bm{x},\,\bm{z},\,\bm{\theta}\right),$ (12)
where $Q_{kk}\left(\bm{g}_{kk};\,\bm{x},\,\bm{z},\,\bm{\theta}\right)$ is defined in (10) while $Q_{kl}\left(\bm{g}_{kl};\,\bm{x},\,\bm{z},\,\bm{\theta}\right)$ is defined by
$Q_{kl}\left(\bm{g}_{kl};\,\bm{x},\,\bm{z},\,\bm{\theta}\right)\;\;\coloneqq\;\;\sum_{i=1}^{N}\sum_{j=1}^{N}g_{ij}\,u_{b}(\bm{x}_{i},\bm{x}_{j},\bm{z}_{i},\bm{z}_{j},\bm{\alpha_{b}},\bm{\beta}_{b})\,z_{ik}\,z_{jl}.$ (13)
In this software development process, each developer team optimally responds to the existing software dependency network. The Nash equilibria of the underlying potential game of the software development process are the networks that maximize the potential function $Q$ (Mele, 2017, 2022; Monderer and Shapley, 1996; Chandrasekhar, 2016; Blume et al., 2013). These are networks such that no developer team is willing to form an additional dependency or delete an existing dependency, when the need arises to update package dependencies. Upon inspecting the stationary distribution $\pi$, it is evident that the Nash equilibria have the highest probability of being observed, because the Nash equilibria maximize $Q$ and hence maximize $\pi$.
Having said that, Nash networks will be inefficient in the presence of externalities. Consider the welfare function defined by the sum of the individual maintainers utilities
$\begin{array}[]{llllll}W(\bm{g},\bm{x},\bm{z},\bm{\theta})&\coloneqq&\displaystyle\sum_{i=1}^{N}U_{i}(\bm{g},\bm{x},\bm{z},\bm{\theta})\vspace{.25cm}\\ &=&Q(\bm{g},\bm{x},\bm{z},\bm{\theta})\,+\,\gamma\,\displaystyle\sum_{k=1}^{K}\sum_{i=1}^{N}\,\sum_{j\neq i}^{N}\,\sum_{r\neq i,j}^{N}g_{ij}g_{ri}z_{ik}z_{jk}z_{rk},\end{array}$ (14)
which is not equal to the potential function $Q(\bm{g},\bm{x},\bm{z},\bm{\theta})$ when $\gamma\neq 0$, so the networks that maximize welfare may not maximize the potential function and are hence not Nash equilibria.
@@ 4 Estimation and Empirical Results
We perform the estimation under the assumption that the observed software dependency network is a sample from the stationary distribution $\pi$ of the discrete-time Markov process described in Section 3.2. This implies that we can estimate the structural parameters based on a single observation of the software dependency network. We sketch a scalable two-step estimation algorithm and refer the interested reader to the appendix for more technical details.
@@ 4.1 Estimation
We introduce a scalable approach to estimating directed network formation models with observable and unobservable types. While scalable estimation methods are available for other models, including stochastic block models (Vu et al., 2013), models which assume that links are independent conditional on latent structure (Mele et al., 2023; Athreya et al., 2018; Kitamura and Laage, 2024), and models for undirected networks (Dahbura et al., 2021; Babkin et al., 2020; Schweinberger and Stewart, 2020), no scalable method for models of directed networks is implemented in publicly available and user-friendly software.
To appreciate the computational challenges involved in estimating directed network formation models, note that the loglikelihood function of the parameter vectors $\bm{\theta}$ and $\bm{\eta}$, given a single observation of a software dependency network $\bm{g}$, is
$\begin{array}[]{lll}\ell(\bm{\theta},\,\bm{\eta})&\coloneqq&\displaystyle\log\sum_{\bm{z}\in\mathcal{Z}}p\left(\bm{z};\,\bm{\eta}\right)\,\pi(\bm{g};\,\bm{x},\,\bm{z},\,\bm{\theta}),\end{array}$ (15)
where $\mathcal{Z}$ is the set of all $K^{N}$ possible configurations of types $\bm{z}$,  $p$ is the multinomial probability of $\bm{z}$ according to (1),  and $\pi$ is the stationary distribution of $\bm{g}$ conditional on observables $\bm{x}$, unobservables $\bm{z}$, and the structural parameter vector $\bm{\theta}$. Maximizing $\ell$ with respect to $\bm{\theta}$ and $\bm{\eta}$ is challenging, because it involves a sum over all $K^{N}$ possible assignments of $N$ packages to $K$ types, which is exponential in $N$. Worse, for each of the $K^{N}$ possible assignments of $N$ packages to $K$ types, the stationary distribution $\pi$ is a function of an intractable normalizing constant, which is the sum of between- and within-type sums, each of which is exponential in the number of pairs of packages $\binom{N_{k}}{2}$ of type $k$ ($k=1,\ldots,K$). Therefore, direct computation of the likelihood function is infeasible unless $N$ is small (e.g., $N\leq 10$).
To develop scalable methods for directed network formation models, we follow Babkin et al. (2020) and decouple the estimation of the unobserved types from the structural utility parameters using a scalable two-step algorithm:
  1. Step 1: Estimate the types of packages.
  2. Step 2: Estimate the structural parameter vector $\bm{\theta}$ conditional on the estimated types.
Step 1
In the first step, we estimate the types of the packages by assuming that there are no externalities ($\gamma=0$) and bounding $\ell$ from below using variational methods, and then using maximization-minorization (MM) algorithm to increase the lower bound on $\ell$. We then assign each package to the type with the highest posterior probability.
We show in Appendix A that the loglikelihood function $\ell$ can be approximated by the following variational lower bound:
$\begin{array}[]{lllll}\ell_{LB}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\,\bm{\Xi})\coloneqq\displaystyle\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{K}\sum_{l=1}^{K}\xi_{ik}\,\xi_{jl}\log\,\pi_{ij,kl}(g_{ij},\bm{x})+\displaystyle\sum_{i=1}^{N}\sum_{k=1}^{K}\xi_{ik}\left(\log\eta_{k}-\log\xi_{ik}\right),\end{array}$ (16)
where $\bm{\Xi}\coloneqq(\xi_{ik})\in[0,1]^{N\times K}$ are auxiliary parameters that serve as approximations of the posterior probabilities that packages $i$ are of type $k$,  $\eta_{k}$ is the probability of being of type $k$,  and $\pi_{ij,kl}$ is the conditional choice probability of a dependency of package $i$ on package $j$, provided $i$ is of type $k$ and $j$ is of type $l$. It is worth noting that $\ell_{LB}$ provides a family of lower bounds on the loglikelihood function $\ell$: Each possible value of the auxiliary parameters $\bm{\Xi}$ provides a lower bound on the loglikelihood function $\ell$. To obtain the best bound on the loglikelihood function $\ell$, we maximize $\ell_{LB}$ with respect to $\bm{\Xi}$.
Maximizing $\ell_{LB}$ with respect to $\bm{\Xi}$ can be cumbersome and time-consuming. Worse, maximization algorithms may be trapped in local maxima of $\ell_{LB}$. Therefore, we construct a surrogate function $M$, which is more convenient to maximize than $\ell_{LB}$:
$M\left(\bm{\Xi};\,\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\Xi}^{(t)}\right)$$\coloneqq$ $\sum_{i=1}^{N}\sum_{j\neq i}^{N}\sum_{k=1}^{K}\sum_{l=1}^{K}\left(\xi_{ik}^{2}\frac{\xi_{jl}^{(t)}}{2\xi_{ik}^{(t)}}+\xi_{jl}^{2}\frac{\xi_{ik}^{(t)}}{2\xi_{jl}^{(t)}}\right)\log\pi_{ij,kl}^{(t)}(g_{ij},\bm{x})$
$+$ $\sum_{i=1}^{N}\sum_{k=1}^{K}\xi_{ik}\left(\log\eta_{k}^{(t)}-\log\xi_{ik}^{(t)}-\frac{\xi_{ik}}{\xi_{ik}^{(t)}}+1\right),$
where $\xi_{ij}^{(t)}$ denotes the estimate of $\xi_{ij}$ at iteration $t$. Relying on the inequality of arithmetic and geometric means and convexity of the negative logarithm, one can show that (4.1) minorizes (16) in the point $\bm{\Xi}^{(t)}$ (see Vu et al., 2013 for further details). Therefore, maximizing (4.1) with respect to $\bm{\Xi}$ forces $\ell_{LB}$ uphill. This step amounts to a quadratic programming problem under box constraints (Stefanov, 2004). This approach along with fast sparse matrix operations, speeds up computations compared with naive variational algorithms that attempt to maximize $\ell_{LB}$ as a function of $\bm{\Xi}$ directly. [^23] See Appendix A for technical details and explanations on how these approximations improve speed of computation. Using these approximations, we can estimate unobserved types at scale while preserving the correlation structure of software dependencies arising from externalities.
Step 2
In the second step, we estimate the structural parameter vector $\bm{\theta}$ by maximizing the pseudolikelihood, which is the product of conditional choice probabilities of links conditional on the types estimated in the first step (Babkin et al., 2020; Martínez Dahbura et al., 2023).
The two-step algorithm sketched above is scalable, in the sense that it scales from small to large networks, and has been tested on undirected networks with a quarter of million units (Dahbura et al., 2021). We provide additional theoretical and computational details in Appendix A and the Online Appendix.